40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 20 16:14:17 2017, gautam, Update, Equipment loan, Impedance test kit borrowed from Downs 
    Reply  Thu Sep 21 01:32:00 2017, gautam, Update, ALS, AUX X Innolight AM measurement running Innolight_AM.pdfInnolight_AM.tar.gzIMG_7599.JPG20170921_203741_TFAG4395A_21-09-2017_115547_FourSquare.pdf
       Reply  Thu Sep 21 01:55:16 2017, rana, Update, ALS, X End table of Shame 20170921_003143.jpg20170921_002430.jpg20170921_002243.jpg20170921_001906.jpg
          Reply  Thu Sep 21 15:23:04 2017, gautam, Omnistructure, ALS, Long cable from LSC->IOO 
          Reply  Fri Sep 29 11:16:52 2017, Steve, Update, ALS, Y End table corrected ETMYf1.jpg
          Reply  Fri Oct 6 17:08:09 2017, Steve, Update, ALS, X End table beam traps corrected c1.jpgc2.jpgc3.jpgc4.jpg
       Reply  Sun Sep 24 20:47:15 2017, rana, Update, Computer Scripts / Programs, RF TF Uncertainties TFAG4395A_21-09-2017_115547_FourSquare.pdfTFAG4395A_21-09-2017_115547.tgz
          Reply  Wed Sep 27 00:20:19 2017, gautam, Update, ALS, More AM sweeps TFAG4395A_26-09-2017_202344_FourSquare.pdflowTemp.tgzTFAG4395A_26-09-2017_231630_FourSquare.pdfhighTemp.tgz
             Reply  Wed Sep 27 23:44:45 2017, gautam, Update, ALS, Proposed PM measurement setup IMG_7609.JPG
Message ID: 13325     Entry time: Thu Sep 21 01:32:00 2017     In reply to: 13324     Reply to this: 13326   13329
Author: gautam 
Type: Update 
Category: ALS 
Subject: AUX X Innolight AM measurement running 

[rana,gautam]

We set up a measurement of the AUX X laser AM today. Some notes:

  • PDA 55 that was installed as a power monitor for the AUX X laser has been moved into the main green beam path - it is just upstream of the green shutter for this measurement.
  • AUX X laser power into the doubling crystal was adjusted by rotating HWP upstream of IR Faraday (original angle was 100, now it is 120), until the DC level of the PDA 55 output was ~2.5V on a scope (high impedance).
  • BNC-T was installed at the PZT input of the Innolight - one arm of the T is terminated to ground via 50 ohms. The purpose of this is to always have the output of the power splitter from the network analyzer RF source drive a 50 ohm load.
  • The output of the Green PDH servo to the Innolight PZT was disconnected downstream of the summing Pomona box - it is now connected to one output of a power splitter (borrowed from SR function generator used to drive the PZT) connected to the RF source output of the AG4395.
  • Other output of power splitter connected to input R of AG4395.
  • PDA55 output has been disconnected from CH5 of the AA board. It is connected to input A of the AG4395 via DC block.

Attachment #1 shows a preliminary scan from tonight - we looked at the region 10kHz-10MHz, with an IF bandwidth of 100Hz, 16 averages, and 801 log-spaced frequencies. The idea was to get an idea of where some promising notches in the AM lie, and do more fine-bandwidth scans around those points. Data + code used to generate this plot in Attachment #2.

Rana points out that some of the AM could also be coming from beam jitter - so to put this hypothesis to test, we will put a lens to focus the spot more tightly onto the PD, repeat the measurement, and see if we get different results.

There were a whole bunch of little illegal things Rana spotted on the EX table which he will make a separate post about.

I am running 40 more scans with the same params for some statistics - should be done by the morning.

Quote:

I borrowed the HP impedance test kit from Rich Abbott today. The purpose is to profile the impedance of the NPRO PZTs, as part of the AUX PDH servo investigations. It is presently at the X-end. I will do the test in the coming days.
 


Update 12:00 21 Sep: Attachment #3 shows schematically the arrangement we use for the AM measurement. A similar sketch for the proposed PM measurement strategy to follow. After lunch, Steve and I will lay out a longish BNC cable from the LSC rack to the IOO rack, from where there is already a long cable running to the X end. This is to facilitate the PM measurement.

Update 18:30 21 Sep: Attachment #4 was generated using Craig's nice plotting utility. The TF magnitude plot was converted to RIN/V by dividing by the DC voltage of the PDA 55 of ~2.3V (assumption is that there isn't significant difference between the DC gain and RF transimpedance gain of the PDA 55 in the measurement band) The right-hand columns are generated by calculating the deviation of individual measurements from the mean value. We're working on improving this utility and aesthetics - specifically use these statistics to compute coherence, this is a work in progress. Git repo details to follow.

There are only 23 measurements (I was aiming for 40) because of some network connectivity issue due to which the script stalled - this is also something to look into. But this sample already suggests that these measurement parameters give consistent results on repeated measurements above 100kHz.

TO CHECK: PDA 55 is in 0dB gain setting, at which it has a BW of 10MHz (claimed in datasheet).


Some math about relation between coherence \gamma_{xy}(f) and standard deviation of transfer function measurements:

\mathrm{SNR}(f) = \sqrt{\frac{\gamma_{xy}^{2}(f)}{1-\gamma_{xy}^{2}(f)}}

\sigma_{xy}^{2} = \frac{1-\gamma_{xy}^{2}(f)}{2N\gamma_{xy}^{2}(f)}|H(f)|^2  --- relation to variance in TF magnitude. We estimate the variance using the usual variance estimator, and can then back out the coherence using this relation.

\sigma_{\theta_{xy}} = \mathrm{tan}^{-1}\left [ \sqrt{\frac{1-\gamma_{xy}^{2}(f)}{2N\gamma_{xy}^{2}(f)}} \right ] --- relation to variance in TF phase. Should give a coherence profile that is consistent with that obtained using the preceeding equation.

It remains to code all of this up into Craig's plotting utility.

Attachment 1: Innolight_AM.pdf  33 kB  Uploaded Thu Sep 21 02:48:22 2017  | Show | Hide all | Show all
Attachment 2: Innolight_AM.tar.gz  13 kB  Uploaded Thu Sep 21 02:49:13 2017
Attachment 3: IMG_7599.JPG  1.202 MB  Uploaded Thu Sep 21 13:12:25 2017  | Hide | Hide all | Show all
IMG_7599.JPG
Attachment 4: 20170921_203741_TFAG4395A_21-09-2017_115547_FourSquare.pdf  730 kB  Uploaded Thu Sep 21 21:39:08 2017  | Hide | Hide all | Show all
20170921_203741_TFAG4395A_21-09-2017_115547_FourSquare.pdf
ELOG V3.1.3-